# The $\infty$-S test via regression quantile affine LASSO

### v1

@@ 1 Introduction
Given a goodness-of-fit measure such as the log-likelihood, the score test calculates the $\ell_{2}$-norm (weighted by the inverse of Fisher information matrix) of the gradient evaluated at the null hypothesis $H_{0}$ and then evaluates whether it is abnormally large compared to a chi-squared distribution. The idea is that if the restricted (under the constraint defined by $H_{0}$) estimator is close the free optimum of the goodness-of-fit measure, the score should not differ from zero significantly. The main advantage of the score test is that it requires only the calculation of the restricted estimate. The score test is also equivalent to the Lagrange multiplier test (Silvey, 1959), and to the F-test (Kohler, 1982) for Gaussian linear regression which assumes the data $(X,\boldsymbol{y})$ are related according to
@@
To derive a pivotal statistic that depends not only on $(X,\boldsymbol{y})$ but also on $(A,\boldsymbol{b})$, Fisher’s F-test uses the Lagrange multiplier technique to define the statistic
$F=\frac{((A\hat{\boldsymbol{\beta}}-\boldsymbol{b})^{\rm T}[A(X^{\rm T}X)^{-1}A^{\rm T}]^{-1}(A\hat{\boldsymbol{\beta}}-\boldsymbol{b}))/m}{(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})^{\rm T}(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})/(n-p)},$
@@
Letting $O_{d_{1},d_{2}}$ be the zero matrix of dimension $d_{1}\times d_{2}$ and $I_{d_{1}}$ be the identity matrix of dimension $d_{1}\times d_{1}$, these exclusion hypotheses correspond to the general hypotheses with $A=[O_{p_{2}\times p_{1}},I_{p_{2}}]$, $\boldsymbol{b}=\boldsymbol{0}_{p_{2}}$ and $m=p_{2}$. In the spirit of the F-test, they derive a statistic
$T_{n}=S_{n}^{\rm T}Q_{n}^{-1}S_{n}/A^{2}(\phi)\rightarrow_{d}\chi^{2}_{p_{2}},$ (4)
where $S_{n}$ is the score under $H_{0}$; see Gutenbrunner et al. (1993, eq. (2.9)) for details. The test statistic is the weighted $\ell_{2}$-norm of the score, and the rescaling by $Q_{n}^{-1}$ makes the statistic asymptotically pivotal and $\chi^{2}$ distributed. In practice $Q_{n}$ must be estimated which entails estimating the sparsity function $s(\tau)=(f(F^{-1}(\tau)))^{-1}$, where $f$ and $F$ are the unknown density and c.d.f. of the errors, respectively; this additional nuisance parameter prevents the test statistic to be close to the asymptotic $\chi^{2}$ distribution even for large sample size $n$ and makes the implementation unstable, as we observed in our empirical experiments in Section 5.
@@
Like the F-test, we propose to derive instead a statistic that directly tackles the desired null hypotheses (2). Our derivation of the $\infty$-S test is based on the least absolute deviation version of the affine LASSO point estimator (Sardy et al., 2022). Affine LASSO is a regularized estimator which regularization parameter can be seen as a Lagrange multiplier to constrain the null hypothesis. So the $\infty$-S test is both a score test and a Lagrange multiplier test.
The article is organized as follows. In Section 2, we derive the $\infty$-S test by defining the affine LAD LASSO (least absolute deviation least absolute shrinkage and selection operator), deriving its zero-thresholding function, and defining the sign-based test statistic that is asymptotically pivotal with respect the nuisance parameters. In Section 3, we discuss optimization issues related to LAD regression with linear constraints and affine LAD LASSO. In Section 4, we extend the $\infty$-S test to testing a quantile of the distribution through the definition of the quantile regression affine LASSO estimator. In Section 5, we apply the $\infty$-S test to four settings: two simple settings to retrieve existing sign tests, a quantile total variation sign test to test for a jump in the tail of time series, and an empirical robustness analysis to non-Gaussianity of the $\infty$-S test and the F-test to match the desired nominal level of the test. The proofs are postponed to an Appendix.
@@ 2 The $\infty$-S test
The test statistic $S$ involves the score of the LAD loss under $H_{0}$ like the statistic $T_{n}$ of Gutenbrunner et al. (1993) with two noticable differences: the test statistic $S$ uses the infinite norm instead of the weighted $\ell_{2}$-norm, and its distribution does not require estimating the sparsity function of Gutenbrunner et al. (1993). The sparisty function is only needed to make the statistic asymptotically $\chi^{2}$under $H_{0}$, not to make the statistic pivotal. Indeed, as Theorem 2 shows, the score is already asymptotically pivotal. Their is an important change of paradigm between the generalization of the sign test by Gutenbrunner et al. (1993) and our new $\infty$-S test: their test has an asymptotic listed distribution (e.g., $F$ or $\chi^{2}$), while our test has an unlisted distribution. Nevertheless, with today’s computer, one can easily sample the distribution of the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$ by Monte Carlo and estimate the test critical value as the empirical $(1-\alpha)$-quantile of $S$.
@@
Consider the linear model (1). To test (2) at the level $\alpha\in(0,1)$, let $c_{\alpha}=F_{S}^{-1}(1-\alpha)$, where $S=\lambda_{0}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}$ is given in (7). The $\infty$-S test’s test function is defined by
@@ 3 Optimization
So the $\infty$-S test can straightforwardly be implemented with current softwares, in particular using the quantreg library in R (Koenker, 2024).
@@ 4 The quantile $\infty$-Sτ test
To test the null hypothesis against the alternative hypothesis (2), in the spirit of quantile regression, one derives the quantile $\infty$-S test to not necessarily fit the data to the median (that is $\tau=1/2$), but to any $\tau$-quantile for some fixed $\tau\in(0,1)$. To that aim, one defines the affine $\rho_{\tau}$-LASSO estimator as a solution to
@@
Given $\tau\in(0,1)$, consider the linear model (9) for the quantiles. To test (2) at the level $\alpha\in(0,1)$, let $c_{\alpha}^{\tau}=F_{S^{\tau}}^{-1}(1-\alpha)$, where $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}^{\tau}$ is given in (14). The quantile $\infty$-Sτ test’s test function is defined by
@@
Based on Theorem 6, the level of the $\infty$-Sτ test is asymptotically the nominal level $\alpha$. Weighting the sign $\boldsymbol{\omega}^{\tau}$ of the residuals by the ranks of the weighted residuals $|\boldsymbol{r}|\cdot\rho_{\tau}(\boldsymbol{r})$ leads to the $\infty$-rankSτ test that has more power than the $\infty$-Sτ test.
@@ 5 Application
Our simulations and application are done with the R software and rely on the quantreg package. We use the rq function to calculate the dual variables $\boldsymbol{\omega}$ necessary to calculate the zero-thresholding function. The critical values are estimated by Monte Carlo simulation with $10^{4}$ runs in all your applications of the $\infty$-S test. The levels and power curves of the tests considered are estimated based on $10^{4}$ simulated data sets. We use the anova.rq function to perform the sign test with asymptotic $\chi^{2}$ distribution whenever it returned a p-value; otherwise the results of the $\chi^{2}$ sign test cannot be reported.
@@ 5.1 The paired $\infty$-S test
Consider two paired measurements $\{(a_{i},b_{i})\}_{i=1,\ldots,n}$, a sample from the model
$A_{i}=\mu_{i}+\tilde{\epsilon}_{i}\quad{\rm and}\quad B_{i}=\mu_{i}+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
@@
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X\beta+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}={\boldsymbol{b}}-{\boldsymbol{a}}$, $X={\boldsymbol{1}}_{n}$, $\beta=\delta$, ${\boldsymbol{\epsilon}}={\tilde{\boldsymbol{\epsilon}}}-{\breve{\boldsymbol{\epsilon}}}$, $A=1$ and $b=0$. Then the test-statistic value is $s=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{y}})|=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{b}}-{\boldsymbol{a}})|=|\sum_{i=1}^{n}(1(b_{i}>a_{i})-1(b_{i}<a_{i}))|$. Recall that the sign test’s test-statistic value is $\tilde{s}=\sum_{i=1}^{n}1(b_{i}>a_{i})$. But $\sum_{i=1}^{n}1(b_{i}>a_{i})+\sum_{i=1}^{n}1(b_{i}<a_{i})=n$, so $s=|2\tilde{s}-n|$ and the $\infty$-S test is equivalent to the existing sign test.
@@ 5.2 The unpaired $\infty$-S test
Consider two unpaired measurements $\{a_{i}\}_{i=1,\ldots,n}$ and $\{b_{i}\}_{i=1,\ldots,n}$, a sample from the model
$A_{i}=\mu+\tilde{\epsilon}_{i}\quad{\rm and}\quad B_{i}=\mu+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
@@
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X{\boldsymbol{\beta}}+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}^{\rm T}=({\boldsymbol{a}}^{\rm T},{\boldsymbol{b}}^{\rm T})$,
@@
${\boldsymbol{\beta}}=(\mu,\delta)^{\rm T}$, ${\boldsymbol{\epsilon}}^{\rm T}=({\tilde{\boldsymbol{\epsilon}}}^{\rm T},{\breve{\boldsymbol{\epsilon}}}^{\rm T})$, $A=(0,1)$ and $b=0$. One gets $\hat{\boldsymbol{\beta}}_{H_{0}}=(m_{\boldsymbol{y}},0)^{\rm T}$, and the test-statistic value is $s=\max(0,|\sum_{i=1}^{n}1(b_{i}>m_{\boldsymbol{y}})-\sum_{i=1}^{n}1(b_{i}<m_{\boldsymbol{y}})|)$. This is the sign test for unpaired data, also known as the median test (Brown and Mood, 1951).
Figure 1: Power of the unpaired $\infty$-Sτ test (same as median) and $\infty$-rankSτ test for $\tau=0.5$, the F-test and asymptotically $\chi^{2}$ sign test as a function of the shift $\delta\in[0,2]$ between the two populations. The horizontal dotted line is the nominal level $\alpha=0.05$. Sample size $n=20$ (left plot) and $n=100$ (right plot).
We empirically compare the level and the power of five tests ($\infty$-S test, $\infty$-rankS test, F-test, median-test and asymptotically $\chi^{2}$ sign test) on two Monte Carlo simulations with $n=20$ and $n=100$ and with Student errors with 3 degrees of freedom. Figure 1 shows the result. All tests satisfy the nominal level of $\alpha=0.05$, the F-test looks quite robust to non-Gaussian errors although losing some power. The $\chi^{2}$ sign test seems to overshoot the level on this regression of small dimension, even for $n$ large.

### v2

# The $\infty$-S test via quantile affine LASSO

@@ Abstract
Keywords: LASSO, nonparametric statistical test, rescaling, robustness, total variation.
@@ 1 Introduction
Given a goodness-of-fit measure such as the log-likelihood, the score test calculates the $\ell_{2}$-norm (weighted by the inverse of Fisher information matrix) of the gradient evaluated at the null hypothesis $H_{0}$ and evaluates whether it is abnormally large compared to a chi-squared distribution. The idea is that if the restricted (under the constraint defined by $H_{0}$) estimator is close the free optimum of the goodness-of-fit measure, the score should not differ from zero significantly. The main advantage of the score test is that it requires only the calculation of the restricted estimate. The score test is equivalent to the Lagrange multiplier test (Silvey, 1959), and to the F-test (Kohler, 1982) for Gaussian linear regression which assumes the data $(X,\boldsymbol{y})$ are related according to
@@
We assume that ${\rm rank}(X)=\min(n,p)$ and ${\rm rank}(X)-m<n$, for instance, the situations where $p<n$ or where $p\geq n$ and $m=p$. To derive a pivotal statistic that depends not only on $(X,\boldsymbol{y})$ but also on $(A,\boldsymbol{b})$, Fisher’s F-test uses the Lagrange multiplier technique to define the statistic
$F=\frac{(A\hat{\boldsymbol{\beta}}-\boldsymbol{b})^{\rm T}[A(X^{\rm T}X)^{-1}A^{\rm T}]^{-1}(A\hat{\boldsymbol{\beta}}-\boldsymbol{b})/m}{(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})^{\rm T}(\boldsymbol{Y}-X\hat{\boldsymbol{\beta}})/(n-p)},$
@@
Letting $O_{d_{1},d_{2}}$ be the zero matrix of dimension $d_{1}\times d_{2}$ and $I_{d_{1}}$ be the identity matrix of dimension $d_{1}\times d_{1}$, these exclusion hypotheses correspond to the general hypotheses (2) with $A=[O_{p_{2}\times p_{1}},I_{p_{2}}]$, $\boldsymbol{b}=\boldsymbol{0}_{p_{2}}$ and $m=p_{2}$. In the spirit of the F-test, they derive a statistic
$T_{n}=S_{n}^{\rm T}Q_{n}^{-1}S_{n}/C_{\varphi}\rightarrow_{d}\chi^{2}_{p_{2}},$ (4)
where $S_{n}$ is the score under $H_{0}$ and $C_{\varphi}$ is some constant; see Gutenbrunner et al. (1993, eq. (2.9)) for details. Their test statistic is the weighted $\ell_{2}$-norm of the score, and the rescaling by $Q_{n}^{-1}$ makes the statistic asymptotically pivotal and $\chi^{2}$ distributed. For that reason we call their test the $\chi^{2}$ sign test. In practice $Q_{n}$ must be estimated which entails estimating the sparsity function $s(\tau)=(f(F^{-1}(\tau)))^{-1}$, where $f$ and $F$ are the unknown density and c.d.f. of the errors, respectively, and $\tau\in(0,1)$ is the quantile of interest; this additional nuisance parameter prevents the test statistic to be close to the asymptotic $\chi^{2}$ distribution even for large sample size $n$ and makes the implementation unstable, as we observe in our empirical experiments in Section 5.
@@
Like the F-test, we derive instead a statistic that directly tackles the desired null hypotheses (2), but as opposed to the F-test, we do not seek an asymptotic $\chi^{2}$ distribution. Our derivation of the $\infty$-S test is based on the least absolute deviation version of the affine LASSO point estimator (Sardy et al., 2022). Affine LASSO is a regularized estimator which regularization parameter can be seen as a Lagrange multiplier to constrain the null hypothesis. So the $\infty$-S test is both a score test and a Lagrange multiplier test.
The article is organized as follows. In Section 2, we derive the $\infty$-S test by defining the affine LAD LASSO (least absolute deviation least absolute shrinkage and selection operator), deriving its zero-thresholding function, and defining the sign-based test statistic that is asymptotically pivotal with respect the nuisance parameters. In Section 3, we discuss optimization issues related to LAD regression with linear constraints and affine LAD LASSO. In Section 4, we extend the $\infty$-S test to testing a quantile of the distribution through the definition of the quantile affine LASSO estimator. In Section 5, we apply the $\infty$-S test to four settings: two simple settings to retrieve existing sign tests, a quantile total variation sign test to test for a jump in the tail of time series, and an empirical robustness analysis to non-Gaussianity of the $\infty$-S test and the F-test to match the desired nominal level of the test. The proofs are postponed to an Appendix.
@@ 2 The $\infty$-S test
The test statistic $S$ involves the score of the LAD loss under $H_{0}$ like the statistic $T_{n}$ of Gutenbrunner et al. (1993) with two noticable differences: the test statistic $S$ uses the infinite norm instead of the weighted $\ell_{2}$-norm, and its distribution does not require estimating the sparsity function of Gutenbrunner et al. (1993). The sparisty function is only needed to make the statistic asymptotically $\chi^{2}$under $H_{0}$, not to make the statistic pivotal. Indeed, as Theorem 2 shows, the score is already asymptotically pivotal. There is an important change of paradigm between the generalization of the sign test by Gutenbrunner et al. (1993) and our new $\infty$-S test: their test has an asymptotic listed distribution (e.g., $F$ or $\chi^{2}$), while our test has an unlisted distribution. Nevertheless, with today’s computer, one can easily sample the distribution of the test statistic $S=\lambda_{0}(\boldsymbol{Y},X,A)$ by Monte Carlo and estimate the test critical value as the empirical $(1-\alpha)$-quantile of $S$. The p-value is also estimated as the percentage of Monte Carlo simulated random variables larger than $\lambda_{0}(\boldsymbol{y},X,A)$.
@@
Consider the linear model (1). To test (2) at a prescribed level $\alpha\in(0,1)$, let $c_{\alpha}=F_{S}^{-1}(1-\alpha)$, where $S=\lambda_{0}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}$ is given in (7). The $\infty$-S test function is defined by
@@
To completely define the test, the rescaling of $(A,\boldsymbol{b})$ must be performed to achieve homogeneous power, that it, to not favor certain alternative hypotheses to the detriment of others. Indeed, the matrices $A$ and $X$ create heteroschedasticity in the random vector $\boldsymbol{W}$ in the test statistics $S=\|\boldsymbol{W}\|_{\infty}$ with $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ before taking the maximum of all entries. Consequently, under $H_{1}$, the entries of $\boldsymbol{W}$ with small deviations will be hidden behind the entries with large deviations. To be democratic between potential entries of $A\boldsymbol{\beta}-\boldsymbol{b}$ not being equal to zero, we propose the following rescaling.
Definition 2 (Homopower rescaling).
Consider the $\infty$-S test with test statistic $S=\|\boldsymbol{W}\|_{\infty}$, where $\boldsymbol{W}=|(AA^{\rm T})^{-1}AX^{\rm T}{\boldsymbol{\omega}}|$ and ${\boldsymbol{\omega}}$ are the sign of the residuals of the constrained LAD under $H_{0}$. Let $F_{W_{k}}$ be the marginal cumulative distribution function of the $k$th entry of $\boldsymbol{W}$ for $k=1,\ldots,m$. Letting $D_{(A,X,{\bf b},\alpha)}$ be the diagonal matrix with $k$th diagonal entry equal to $d_{k}:=F_{W_{k}}^{-1}(1-\alpha)$, homopower rescaling rescales $A$ to $D_{(X,A,{\bf b},\alpha)}A$ and $\boldsymbol{b}$ to $D_{(X,A,{\bf b},\alpha)}\boldsymbol{b}$ before applying the test.
@@ 3 Optimization
So the $\infty$-S test can straightforwardly be implemented with current softwares, in particular using the quantreg package (Koenker, 2024) in R.
@@ 4 The quantile $\infty$-Sτ test
To test the null hypothesis against the alternative hypothesis (2), in the spirit of quantile regression, one derives the quantile $\infty$-S test to not necessarily fit the data to the median (that is $\tau=1/2$), but to any $\tau$-quantile for some fixed $\tau\in(0,1)$. To that aim, we define affine $\rho_{\tau}$-LASSO as a solution to
@@
Given $\tau\in(0,1)$, consider the linear model (9) for the quantiles. To test (2) at a prescribed level $\alpha\in(0,1)$, let $c_{\alpha}^{\tau}=F_{S^{\tau}}^{-1}(1-\alpha)$, where $S^{\tau}=\lambda_{0}^{\tau}(\boldsymbol{Y},X,A)$ and $\boldsymbol{Y}$ is the response random vector for the linear model under $H_{0}$ and the function $\lambda_{0}^{\tau}$ is given in (14). The quantile $\infty$-Sτ test function is defined by
@@
Based on Theorem 6, the level of the $\infty$-Sτ test is asymptotically the nominal level $\alpha$. Weighting the sign $\boldsymbol{\omega}^{\tau}$ of the residuals by the ranks of the weighted residuals $|\boldsymbol{r}|\cdot\rho_{\tau}(\boldsymbol{r})$ leads to the $\infty$-rankSτ test.
@@ 5 Application
Our simulations and application are done with the R software and rely on the quantreg package (Koenker, 2024). We use the rq function to calculate the dual variables $\boldsymbol{\omega}$ necessary to calculate the zero-thresholding function. The critical values are estimated by Monte Carlo simulation with $10^{4}$ runs in all our applications of the $\infty$-S test. The levels and power curves of the tests considered are estimated based on $10^{4}$ simulated data sets. We use the anova.rq function to perform the $\chi^{2}$ sign test whenever it returned a p-value; otherwise the results of the $\chi^{2}$ sign test cannot be reported.
@@ 5.1 The paired $\infty$-S test
Consider two paired measurements $\{(u_{i},v_{i})\}_{i=1,\ldots,n}$, a sample from the model
$U_{i}=\mu_{i}+\tilde{\epsilon}_{i}\quad{\rm and}\quad V_{i}=\mu_{i}+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
@@
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X\beta+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}={\boldsymbol{v}}-{\boldsymbol{u}}$, $X={\boldsymbol{1}}_{n}$, $\beta=\delta$, ${\boldsymbol{\epsilon}}={\tilde{\boldsymbol{\epsilon}}}-{\breve{\boldsymbol{\epsilon}}}$, $A=1$ and $b=0$. Then the test-statistic value is $s=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{y}})|=|\sum_{i=1}^{n}{\rm sign}({\boldsymbol{v}}-{\boldsymbol{u}})|=|\sum_{i=1}^{n}(1(v_{i}>u_{i})-1(v_{i}<u_{i}))|$. Recall that the sign test’s statistic value is $\tilde{s}=\sum_{i=1}^{n}1(v_{i}>u_{i})$. But $\sum_{i=1}^{n}1(v_{i}>u_{i})+\sum_{i=1}^{n}1(v_{i}<u_{i})=n$, so $s=|2\tilde{s}-n|$ and the $\infty$-S test is equivalent to the existing sign test.
@@ 5.2 The unpaired $\infty$-S test
Consider two unpaired measurements $\{u_{i}\}_{i=1,\ldots,n}$ and $\{v_{i}\}_{i=1,\ldots,n}$, a sample from the model
$U_{i}=\mu+\tilde{\epsilon}_{i}\quad{\rm and}\quad V_{i}=\mu+\delta+\breve{\epsilon}_{i},\ i=1,\ldots,n,$
@@
with the $\infty$-S test, one writes the model as ${\boldsymbol{y}}=X{\boldsymbol{\beta}}+{\boldsymbol{\epsilon}}$ with ${\boldsymbol{y}}^{\rm T}=({\boldsymbol{u}}^{\rm T},{\boldsymbol{v}}^{\rm T})$,
@@
${\boldsymbol{\beta}}=(\mu,\delta)^{\rm T}$, ${\boldsymbol{\epsilon}}^{\rm T}=({\tilde{\boldsymbol{\epsilon}}}^{\rm T},{\breve{\boldsymbol{\epsilon}}}^{\rm T})$, $A=(0,1)$ and $b=0$. One gets $\hat{\boldsymbol{\beta}}_{H_{0}}=(m_{\boldsymbol{y}},0)^{\rm T}$, and the test-statistic value is $s=\max(0,|\sum_{i=1}^{n}1(v_{i}>m_{\boldsymbol{y}})-\sum_{i=1}^{n}1(v_{i}<m_{\boldsymbol{y}})|)$. This is the sign test for unpaired data, also known as the median test (Brown and Mood, 1951).
Figure 1: Power of the unpaired $\infty$-Sτ test (same as median test) and $\infty$-rankSτ test for $\tau=0.5$, the F-test and asymptotically $\chi^{2}$ sign test as a function of the shift $\delta\in[0,2]$ between the two populations. The horizontal dotted line is the nominal level $\alpha=0.05$. Sample size $n=20$ (left plot) and $n=100$ (right plot).
We compare the level and the power of four tests ($\infty$-S or median test, $\infty$-rankS test, F-test and $\chi^{2}$ sign test) on two Monte Carlo simulations with $n=20$ and $n=100$ and with Student errors with 3 degrees of freedom. Figure 1 shows that all tests satisfy the nominal level of $\alpha=0.05$ and the F-test looks quite robust to non-Gaussian errors although losing some power. The $\chi^{2}$ sign test seems to overshoot the level, even for $n$ large, however.
5.3 Homopower rescaling
We perform a toy simulation to outline the importance of homopower rescaling (see Definition 2) the two components $A$ and $\boldsymbol{b}$ of the linear null hypothesis (2). Here $n=100$, $p=20$, $m=2$, $\tau=0.5$, $\boldsymbol{b}=\boldsymbol{0}_{m}$ and $A=[C,O_{m\times(p-m)}]$ with
$C=\left(\begin{array}[]{ccc}3&0\\ 0&1\end{array}\right).$
The regression matrix $X$ has entries sampled from a standard Gaussian random variable, the null model is $H_{0}:\ \boldsymbol{\beta}=\boldsymbol{0}$, and we consider two alternative hypotheses
$H_{1,1}:\ \boldsymbol{\beta}=\delta\boldsymbol{e}_{1}\quad{\rm and}\quad H_{1,2}:\ \boldsymbol{\beta}=\delta\boldsymbol{e}_{2},$
where $\boldsymbol{e}_{i}$ is the $i$th canonical vector. With the unrescaled $\infty$-S test, one expects the power to be low under $H_{1,1}$ and good under $H_{1,2}$, while with the rescaled $\infty$-S test, one expects the power to be high under both $H_{1,1}$ and $H_{1,2}$. Figure 2 illustrates the improvement thanks to homopower rescaling. Not only the $\infty$-S test, but many tests would profit from this rescaling in terms of power. This example is certainly a toy example, but what we observe here also occurs naturally owing to the relation between $A$ and $X$ that is not controlled by the user, except possibly in ANOVA. So homopower rescaling should be employed.
Figure 2: Homopower rescaling. Top: boxplots of Monte Carlo simulated bivariate ($m=2$) random vector $W$ unrescaled (left) and rescaled (right). Bottom: power plots of the unrescaled (red) and rescaled (black) $\infty$-S test under $H_{1,1}$ (left) and $H_{1,2}$ (right).